# arXiv:2603.00183v1[cs.SE] 26 Feb 2026

### v1

@@ Abstract
Methods: We conduct a snowballing review (SR) on TCP, implement a comprehensive platform for TCP research (TCPFramework), analyze existing evaluation metrics and propose two new ones ($r\text{APFD}_{C}$ and ATR), and develop a family of ensemble TCP methods called approach combinators.
@@
keywords:
test case prioritization , regression test prioritization , continuous integration , ensemble methods , code representation
journal: Journal of Systems and Software
[…]
[inst1]organization=Wroclaw University of Science and Technology, city=Wroclaw, country=Poland
{graphicalabstract}[Uncaptioned image]
{highlights}
A new, wide snowballing review of TCP following modern review guidelines.
Implementation of an open-source, unified TCP research infrastructure (TCPFramework).
Analysis of existing TCP evaluation metrics and a proposition of two new ones.
Introduction of TCP approach combinators: mixers, interpolators, and tiebreakers.
Empirical evaluation of proposed methods that achieve results comparable to SOTA.
@@ 1 Introduction
Software bugs can frustrate users and reduce their satisfaction Saha et al. [2014] or even cause health and safety risks in certain areas. However, bugs can occur in all software systems Ajorloo et al. [2024]. To ensure that a program behaves as expected, software testing is performed. In particular, regression testing (RT) is commonly used to ensure that changes to the system did not introduce unwanted behavior Washizaki [2024].
Due to the volume of executed tests, RT can be costly, both in terms of time and money, with estimates stating that it may comprise between 33% and 50% of software expenses Washizaki [2024], Myers et al. [2011], Khatibsyarbini et al. [2018]. Regression testing is at the same time an important part of Agile and continuous integration (CI), where automated RT suites are run on each build Washizaki [2024], Haghighatkhah et al. [2018], Prado Lima and Vergilio [2020], Marijan [2023]. However, CI processes often require rapid turnaround, to avoid delays in build cycles, and to provide software developers with the necessary feedback Bagherzadeh et al. [2022], Marijan [2023].
Common approaches to speed up the RT process include test suite minimization, test case selection, and test case prioritization Washizaki [2024], as systematized by Yoo and Harman [2012]. In particular, test case prioritization (TCP), also called regression test prioritization (RTP), aims to reorder test cases to produce the permutation that would detect possible faults as early as possible. It allows all test cases to run in the worst case, but also permits early termination, for example, when the first failing case is encountered Yoo and Harman [2012].
As further discussed in Section˜2, over the years, hundreds of TCP methods have been researched. The technique is also used by practitioners, finding success in large software development entities such as Google Memon et al. [2017] and Microsoft Czerwonka et al. [2011]. However, there are still opportunities for improvement.
@@
  • 1.
    A new snowballing literature review on the topic of test case prioritization that follows modern review guidelines Kitchenham et al. [2023], Wohlin [2014] and is among the largest reviews of the TCP literature in terms of the count of identified works.
  • 2.
@@ 2 Related work
In order to summarize the current state of the art of TCP, and to locate opportunities for further improvement, a snowballing review (SR) was undertaken. This section describes it in terms of the planning phase, the search procedure, and its results. Then, the most relevant related work is outlined, and an analysis of available TCP evaluation metrics is performed.
@@ 2.1 Literature review process
Systematic review is a technique to systematically discover, evaluate, and synthesize scientific research related to a specific topic. Compared to traditional unsystematic reviews, systematic reviews are conducted formally, using a set of methodological steps and approaches listed in a review protocol Kitchenham et al. [2007]. The following review adheres to the SEGRESS guidelines Kitchenham et al. [2023] and uses the snowballing method introduced by Wohlin [2014].
@@
Upon formulating the research objectives, the review protocol was written. According to the available recommendations, the protocol should, among others, describe Kitchenham et al. [2007]: the motivation for the SR, the previously defined research questions, search methods (including study selection criteria), data extraction, and analysis strategies. The protocol attempts to adhere to the PRISMA-P statement Moher et al. [2015], as SEGRESS Kitchenham et al. [2023] recommends. It should be noted that the protocol was written before the subsequent SR steps were carried out to reduce bias and published online as Section 1 of the online Appendix to this paper [^3] https://madeyski.e-informatyka.pl/download/ChojnackiMadeyski25Appendix.pdf.
@@
The search process used in the SR follows the snowballing approach recommended by Wohlin [2014]. The iterative nature of this technique provides stable results. Each encountered article is evaluated using the inclusion and exclusion criteria noted in the review protocol linked above.
In order to adapt the search process to the circumstances of the research, a few modifications, justified below, were introduced. It should be noted that these changes always influence only the recommendations, never breaking any hard rules established by the guidelines. First, since we previously gathered a list of unmethodically selected publications closely related to the studied research area, it is used to form the snowballing start set. Secondly, the author search, consisting of directly contacting the researchers associated with identified papers, is infeasible to carry out in the research timeframe. Alternatively, all works by the authors appearing at least five times in the final list are to be analyzed manually without directly asking them.
Google Scholar [^4] https://scholar.google.com is used for the forward snowballing and author search, the Scopus [^5] https://www.scopus.com/search/form.uri database is used to construct the start set, and the reference list contained in the articles themselves is used for backward snowballing. The proposed modifications aim to improve the reliability of the SR and are caused by practical limitations. The altered search process is shown on Figure˜1. Compared to the original process presented by Wohlin [2014], the modified components are colored red, while the newly introduced component is colored green.
Figure 1: Stages of the altered search procedure.
The list of initial papers contains the works of Yang et al. [2023], Marijan [2023], Yaraghi et al. [2023], and Bagherzadeh et al. [2022]. Their keywords were analyzed and used to construct the Scopus search query shown below in Figure˜2. The test case prioritization and test prioritization were the most common keywords across initial papers and the SUBJAREA, PUBSTAGE and LANGUAGE filters come directly from the SR inclusion and exclusion criteria. The somewhat arbitrary publication year constraint was introduced to target only recent studies (from the past five years), while also including all initial papers. The search returned a total of 440 documents. Most notably, all initial papers were also returned by the query. This ensures that the filters were not too restrictive.
@@
The result list was then manually traversed, using inclusion and exclusion criteria, to identify works that could act as the start set. The guidelines of Wohlin [2014] state that a good start set should include papers from different communities, publishers, years, and authors. In the end, the following articles were selected: Bagherzadeh et al. [2022], Huang et al. [2020], Vescan et al. [2024], Mukherjee and Patnaik [2021]. It should be noted that the first entry was previously included in the initial paper set, which is permitted by the SR protocol. Thus, while the start set includes four papers, only three new papers were introduced in this stage. The constructed set spans multiple years, there are no common authors between any two papers in the list, the identified authors come from different communities, and the articles were all published in distinct venues. The selected works contain a total of 274 references and 222 citations (duplicates included), which should provide sufficient resources for the first iteration of snowballing. Subsequently, multiple iterations of snowballing were conducted.
2.2 Snowballing literature review results
@@
The identified works were spread over multiple decades, with the first published in 1999 by Rothermel et al. [1999] and the last in 2025. The methods used for the SR should ensure that it is unbiased in terms of publication years. The total number of articles from different years is shown in Figure˜3.
@@
A total of 54 articles introduced new datasets, 267 articles presented novel algorithms, 32 articles were secondary studies, and 28 articles were dedicated to empirically comparing existing solutions. Some of the identified surveys cover narrower or wider areas than the proposed SR. Numerous also did not disclose the number of primary studies analyzed. Since the discussed works are not suitable for a comparison with the proposed SR, of the 32 identified secondary studies, only eight were handpicked and chronologically presented in Table˜2.
@@
2012 Singh et al. [2012] 1997 – 2011 65
2013 Catal and Mishra [2013] 2001 – 2011 120
2017 de S. Campos Junior et al. [2017] 1999 – 2016 90
2018 Khatibsyarbini et al. [2018] 1999 – 2016 80
2020 de Castro-Cabrera et al. [2020] 2017 – 2019 320
2020 Prado Lima and Vergilio [2020] 2009 – 2019 35
2021 Mukherjee and Patnaik [2021] 2001 – 2018 90
2023 Abubakar et al. [2023] 2002 – 2021 250
@@
In general, the thoroughness of the proposed review is difficult to compare with other studies, as they contained diverse research questions and years covered, and it is possible that some other reviews were missed. However, it can be said that the conducted study is among the largest SRs of TCP in terms of the number of identified publications.
@@
The accumulated data was used to answer SRRQ1. Since TCP can be applied to almost any software product where testing is used, nearly any software repository can be adapted to act as a dataset. Some solutions followed this approach, with the most common choice being Software-artifact Infrastructure Repository projects [^6] https://sir.csc.ncsu.edu (SIR): Unix utilities (i.e., bash, grep, sed), Siemens programs, JMeter, JTopas, Ant Prado Lima and Vergilio [2020], Singh et al. [2012], Catal and Mishra [2013], de S. Campos Junior et al. [2017]. Other identified subject programs include compilers (GCC, javac, Jikes), JDepend, Gradebook, Tomcat, Camel, and Drupal.
@@
Some solutions instead opt for processed test suite datasets, such as the Google Shared Dataset of Test Suite Results [^7] https://code.google.com/archive/p/google-shared-dataset-of-test-suite-results (GSDTSR), Defects4j [^8] https://github.com/rjust/defects4j, and Paint Control [^9] https://bitbucket.org/HelgeS/retecs/src/master/DATA Prado Lima and Vergilio [2020], de Castro-Cabrera et al. [2020]. The RTPTorrent dataset is the only identified data source supported by a dedicated research paper Mattis et al. [2020].
@@
In less complicated problem statements, TCP can be applied to simple fault matrices, which state whether a given test case can detect a given fault. An example of such a dataset is the BigFaultMatrix [^10] https://github.com/dathpo/test-case-prioritisation-ga. Finally, synthetically introduced faults (mutants) might be a viable evaluation method for TCP algorithms. Do and Rothermel [2006] conclude that mutants can be used as a low-cost dataset construction method. However, Luo et al. [2018] highlight that TCP algorithms that perform the best on sets of mutants might perform worse on real-world data.
In general, the answer to SRRQ1 is that there are many available TCP evaluation datasets and subject programs; however, none of them are universally used. The choice of the dataset should be influenced by the exact problem statement and other factors.
Subsequently, the results of the study were used to answer SRRQ2. Here, since TCP has many subdomains, each with its own limitations with regard to data availability, time constraints, and more, it is difficult to directly compare two TCP algorithms. Singh et al. [2012] noted that more empirical comparisons of existing techniques should be performed. However, to directly compare two methods, they must work on the same definitions, metrics, and datasets. Catal and Mishra [2013] offer similar recommendations. Furthermore, the authors state the need for more empirical evaluations of the proposed techniques, noting that a quarter of the identified TCP methods did not use any evaluation metrics to assess effectiveness. Although the cited works are not recent, their findings unfortunately still apply.
@@
Multiple studies attempted to compare existing TCP methods. In 2018, Haghighatkhah et al. [2018] analyzed diversity-based and history-based test prioritization applied to six subject programs. Luo et al. [2016] instead compared static and dynamic test case prioritization techniques. Marijan [2023] compared ML approaches with heuristic-based techniques, finding that LambdaRank with neural network and the ROCKET Marijan et al. [2013] algorithm are the most effective among the tested. The ROCKET technique is also the second most widely used baseline in the evaluation of TCP, after random test case ordering, according to Prado Lima and Vergilio [2020].

### v2

# Test Case Prioritization: A Systematic Review using Snowballing and TCPFramework with Approach Combinators

@@ Abstract
Methods: We conduct a systematic review (SR) using snowballing on TCP, implement a comprehensive platform for TCP research (TCPFramework), analyze existing evaluation metrics and propose two new ones ($r\text{APFD}_{C}$ and ATR), and develop a family of ensemble TCP methods called approach combinators.
@@
Keywords: test case prioritization; regression test prioritization; continuous integration; ensemble methods; code representation
@@ 1 Introduction
Software bugs can frustrate users and reduce their satisfaction (Saha et al., 2014) or even cause health and safety risks in certain areas. However, bugs can occur in all software systems (Ajorloo et al., 2024). To ensure that a program behaves as expected, software testing is performed. In particular, regression testing (RT) is commonly used to ensure that changes to the system did not introduce unwanted behavior (Washizaki, 2024).
Due to the volume of executed tests, RT can be costly, both in terms of time and money, with estimates stating that it may comprise between 33% and 50% of software expenses (Washizaki, 2024; Myers et al., 2011; Khatibsyarbini et al., 2018). Regression testing is at the same time an important part of Agile and continuous integration (CI), where automated RT suites are run on each build (Washizaki, 2024; Haghighatkhah et al., 2018; Prado Lima and Vergilio, 2020; Marijan, 2023). However, CI processes often require rapid turnaround, to avoid delays in build cycles, and to provide software developers with the necessary feedback (Bagherzadeh et al., 2022; Marijan, 2023).
Common approaches to speed up the RT process include test suite minimization, test case selection, and test case prioritization (Washizaki, 2024), as systematized by Yoo and Harman (2012). In particular, test case prioritization (TCP), also called regression test prioritization (RTP), aims to reorder test cases to produce the permutation that would detect possible faults as early as possible. It allows all test cases to run in the worst case, but also permits early termination, for example, when the first failing case is encountered (Yoo and Harman, 2012).
As further discussed in Section˜2, over the years, hundreds of TCP methods have been researched. The technique is also used by practitioners, finding success in large software development entities such as Google (Memon et al., 2017) and Microsoft (Czerwonka et al., 2011). However, there are still opportunities for improvement.
@@
  • A new systematic review using snowballing on the topic of test case prioritization that follows modern review guidelines (Kitchenham et al., 2023; Wohlin, 2014) and is one of the largest reviews of the TCP literature in terms of the count of identified works.
@@ 2 Related work
In order to summarize the current state of the art of TCP, and to locate opportunities for further improvement, a systematic review (SR) using snowballing was undertaken. This section describes it in terms of the planning phase, the search procedure, and its results. Then, the most relevant related work is outlined, and an analysis of available TCP evaluation metrics is performed.
@@ 2.1 Literature review process
Systematic review is a technique to systematically discover, evaluate, and synthesize scientific research related to a specific topic. Compared to traditional unsystematic reviews, systematic reviews are conducted formally, using a set of methodological steps and approaches listed in a review protocol (Kitchenham et al., 2007). The following review adheres to the SEGRESS guidelines (Kitchenham et al., 2023) and uses the snowballing method introduced by Wohlin (2014).
@@
Upon formulating the research objectives, the review protocol was written. According to the available recommendations, the protocol should, among others, describe (Kitchenham et al., 2007): the motivation for the SR, the previously defined research questions, search methods (including study selection criteria), data extraction, and analysis strategies. The protocol attempts to adhere to the PRISMA-P statement (Moher et al., 2015), as SEGRESS (Kitchenham et al., 2023) recommends. It should be noted that the protocol was written before the subsequent SR steps were carried out to reduce bias and published online as Section 1 of the online Appendix to this paper [^1] https://madeyski.e-informatyka.pl/download/ChojnackiMadeyski25Appendix.pdf.
@@
The search process used in the SR follows the snowballing approach recommended by Wohlin (2014). Each encountered article is evaluated using the inclusion and exclusion criteria noted in the review protocol linked above. Google Scholar [^2] https://scholar.google.com is used for forward snowballing, the Scopus [^3] https://www.scopus.com/search/form.uri database is used to construct the start set, and the reference lists contained in the articles themselves are used for backward snowballing. Figure˜1 illustrates the implemented search process together with a flowchart summarizing the number of records remaining after each stage.
Figure 1: Workflow of the search process and flowchart of paper counts at each stage.
The presented search process differs from the original snowballing guidelines in two minor respects. These changes were implemented to accommodate the study’s time constraints and do not violate any principles specified in the guidelines.
First, Wohlin (2014) recommends directly contacting all researchers associated with the included papers after the snowballing iterations, while also allowing alternative methods to identify additional studies. As contacting authors was infeasible within the research timeframe, we instead analyzed their publication records. Specifically, we systematically examined the Google Scholar pages of authors appearing at least five times in the paper collection to identify additional relevant articles. This threshold serves as a pragmatic heuristic to focus only on recurring contributors to the field and to reduce noise from one-off publications. We refer to this stage as the author search.
Secondly, since we previously gathered a list of unmethodically selected publications closely related to the studied research area, it is used to form the snowballing start set. The list of initial papers contains the works of Yang et al. (2023), Marijan (2023), Yaraghi et al. (2023), and Bagherzadeh et al. (2022). Their keywords were analyzed and used to construct the Scopus search query shown in Figure˜2. The test case prioritization and test prioritization were the most common keywords across initial papers and the SUBJAREA, PUBSTAGE and LANGUAGE filters come directly from the SR inclusion and exclusion criteria. The somewhat arbitrary publication year constraint was introduced to target only recent studies (from the past five years), while also including all initial papers. The search returned a total of 440 documents. Most notably, all initial papers were also returned by the query. This ensures that the filters were not too restrictive.
@@
The result list was then manually traversed, using inclusion and exclusion criteria, to identify works that could act as the start set. The guidelines of Wohlin (2014) state that a good start set should include papers from different communities, publishers, years, and authors. In the end, the following articles were selected: Bagherzadeh et al. (2022), Huang et al. (2020), Vescan et al. (2024), Mukherjee and Patnaik (2021). It should be noted that the first entry was previously included in the initial paper set, which is permitted by the SR protocol. Thus, while the start set includes four papers, only three new papers were introduced in this stage. The constructed set spans multiple years, there are no common authors between any two papers in the list, the identified authors come from different communities, and the articles were all published in distinct venues. The selected works contain a total of 274 references and 222 citations (duplicates included), which should provide sufficient resources for the first iteration of snowballing. Subsequently, multiple iterations of snowballing were conducted.
2.2 Systematic review results
@@
The identified works were spread over multiple decades, with the first published in 1999 by Rothermel et al. (1999) and the last in 2025. The methods used for the SR should ensure that it is unbiased in terms of publication years. The total number of articles from different years is shown in Figure˜3.
@@
A total of 54 articles introduced new datasets, 267 articles presented novel algorithms, 32 articles were secondary studies, and 28 articles were dedicated to empirically comparing existing solutions. Some of the identified surveys cover narrower or wider areas than the proposed SR. Numerous also did not disclose the number of primary studies analyzed. Since the discussed works are not suitable for a comparison with the proposed SR, of the 32 identified secondary studies, only eight were handpicked and chronologically presented in Table˜2. Our study reports 292 works rather than 324, as this figure represents only the analyzed primary studies, whereas the total of 324 identified works also includes secondary studies.
@@
2012 Singh et al. (2012) 1997 – 2011 65
2013 Catal and Mishra (2013) 2001 – 2011 120
2017 de S. Campos Junior et al. (2017) 1999 – 2016 90
2018 Khatibsyarbini et al. (2018) 1999 – 2016 80
2020 de Castro-Cabrera et al. (2020) 2017 – 2019 320
2020 Prado Lima and Vergilio (2020) 2009 – 2019 35
2021 Mukherjee and Patnaik (2021) 2001 – 2018 90
2023 Abubakar et al. (2023) 2002 – 2021 250
@@
In general, the thoroughness of the proposed review is difficult to compare with other studies, as they contained diverse research questions and years covered, and it is possible that some other reviews were missed. Nevertheless, the conducted study represents one of the largest SRs of TCP, in terms of the number of identified publications. Among the analyzed surveys, only de Castro-Cabrera et al. (2020) included a higher number of works, when considering raw counts.
@@
The accumulated data was used to answer SRRQ1. Since TCP can be applied to almost any software product where testing is used, nearly any software repository can be adapted to act as a dataset. Some solutions followed this approach, with the most common choice being Software-artifact Infrastructure Repository projects [^4] https://sir.csc.ncsu.edu (SIR): Unix utilities (i.e., bash, grep, sed), Siemens programs, JMeter, JTopas, Ant (Prado Lima and Vergilio, 2020; Singh et al., 2012; Catal and Mishra, 2013; de S. Campos Junior et al., 2017). Other identified subject programs include compilers (GCC, javac, Jikes), JDepend, Gradebook, Tomcat, Camel, and Drupal.
@@
Some solutions instead opt for processed test suite datasets, such as the Google Shared Dataset of Test Suite Results [^5] https://code.google.com/archive/p/google-shared-dataset-of-test-suite-results (GSDTSR), Defects4j [^6] https://github.com/rjust/defects4j, and Paint Control [^7] https://bitbucket.org/HelgeS/retecs/src/master/DATA (Prado Lima and Vergilio, 2020; de Castro-Cabrera et al., 2020). The RTPTorrent dataset is the only identified data source supported by a dedicated research paper (Mattis et al., 2020).
@@
In less complicated problem statements, TCP can be applied to simple fault matrices, which state whether a given test case can detect a given fault. An example of such a dataset is the BigFaultMatrix [^8] https://github.com/dathpo/test-case-prioritisation-ga. Finally, synthetically introduced faults (mutants) might be a viable evaluation method for TCP algorithms. Do and Rothermel (2006) conclude that mutants can be used as a low-cost dataset construction method. However, Luo et al. (2018) highlight that TCP algorithms that perform the best on sets of mutants might perform worse on real-world data.
In summary, the answer to SRRQ1 reveals a significant fragmentation in the field: while numerous TCP evaluation datasets exist (SIR programs, GSDTSR, Defects4j, Paint Control, BigFaultMatrix, RTPTorrent, and various individual repositories), no universally adopted benchmark has emerged. This lack of standardization makes cross-study comparisons difficult. Among the identified datasets, RTPTorrent stands out as the only one supported by a dedicated methodological paper that explicitly addresses common dataset quality issues (Mattis et al., 2020). For these reasons, we selected RTPTorrent for our empirical evaluation (see Section˜4).
Subsequently, addressing SRRQ2 proved challenging due to the heterogeneity of the field. Our findings align with earlier observations by Singh et al. (2012) and Catal and Mishra (2013): TCP algorithms are rarely compared under identical conditions (the same datasets, metrics, and definitions), and a significant portion of the proposed methods lack empirical evaluation.
@@
Despite these limitations, some patterns emerge from existing comparative studies. In 2018, Haghighatkhah et al. (2018) analyzed diversity-based and history-based test prioritization applied to six subject programs. Luo et al. (2016) instead compared static and dynamic test case prioritization techniques. Most notably, Marijan (2023) compared ML approaches with heuristic-based techniques, finding that ROCKET (Marijan et al., 2013) performs comparably to or better than many ML-based methods. The ROCKET technique is also the second most widely used baseline in the evaluation of TCP, after random test case ordering, according to Prado Lima and Vergilio (2020). Based on these findings, we include ROCKET as a primary baseline in our evaluation.
To summarize, the SR revealed that the TCP research landscape is characterized by methodological fragmentation: diverse datasets, incompatible metrics, and limited cross-study comparability. These findings directly informed our research choices: we selected RTPTorrent as the most rigorously documented dataset, adopted ROCKET as a state-of-the-art heuristic baseline, and designed approach combinators as training-free methods that do not depend on specific historical data formats or require extensive tuning. Table˜3 provides a systematic mapping of how specific SR findings informed key methodological decisions in this work.
Table 3: Mapping of SR findings to research decisions.
Research decision Related SR publications How SR informed the decision
Dataset: RTPTorrent (Mattis et al., 2020) SRRQ1 revealed dataset fragmentation (SIR, GSDTSR, Defects4j, Paint Control, BigFaultMatrix, etc.); RTPTorrent is the only dataset with a dedicated methodological paper addressing data quality issues.
Baseline: ROCKET (Marijan et al., 2013),
(Marijan, 2023),
(Prado Lima and Vergilio, 2020)
SRRQ2: ROCKET is the second most widely used TCP baseline after random ordering; performs comparably to or better than many ML-based methods.
Baseline: DBP (Zhou et al., 2021) Identified through SR as a recent heuristic method meeting inclusion criteria for baseline selection.
Baseline: DFE (Kim and Porter, 2002),
(Mattis et al., 2020)
Used as a non-trivial baseline in RTPTorrent; represents simple exponential-smoothing TCP approaches.
Approach
combinators (RQ1)
(Huang et al., 2021),
(Cao et al., 2022),
(Mondal and Nasre, 2021),
(Hassan et al., 2024)
SR identified a gap: prior attempts at combining TCP methods were limited to ML-based approaches (requiring training) or simple voting schemes; no training-free combinators existed.
Metric: NTR (Prado Lima and Vergilio, 2022),
(Zhao et al., 2023)
SR identified NTR as suitable for evaluating CI applicability; inspired the proposed ATR metric addressing its limitations.
Metric: $r\text{APFD}$ (Zhao et al., 2023) SR identified limitations of APFD (inconsistent value ranges); $r\text{APFD}$ addresses this; inspired the proposed $r\text{APFD}_{C}$ metric.
In the following sections, we present the key theoretical topics relevant to this work, drawing primarily on information obtained from the literature review.